0.1 List of Mathematical Symbols

In the following tables, you find all the symbols normally accessible from math mode.

To use the symbols listed in Tables 12–16,^{1} the package amssymb must be loaded in the preamble of the document and the AMS math fonts must be installed, on the system. If the AMS package and fonts are not installed, on your system, have a look at

CTAN:/tex-archive/macros/latex/packages/amslatex

Table 1: Math Mode Accents.

 â \hat{a}     ǎ \check{a}   ã \tilde{a}     á \acute{a}
 à \grave{a}   ȧ \dot{a}     ä \ddot{a}      ă \breve{a}
 ā \bar{a}     a⃗ \vec{a}     Â \widehat{A}   Ã \widetilde{A}

Table 2: Lowercase Greek Letters.

    α \alpha  θ \theta  o o  υ \upsilon
    β \beta  ϑ \vartheta π \pi  ϕ \phi
    γ \gamma  ι \iota  ϖ \varpi  φ \varphi
    δ \delta  κ \kappa  ρ \rho  χ \chi
    ϵ \epsilon  λ \lambda  ϱ \varrho  ψ \psi
    ε \varepsilon μ \mu  σ \sigma  ω \omega
    ζ \zeta  ν \nu  ς \varsigma
    η \eta  ξ \xi  τ \tau

1These tables were derived from symbols.tex by David Carlisle and subsequently changed extensively as suggested by Josef Tkadlec.

\underline{2}

Table 3: Uppercase Greek Letters.

                 Γ \Gamma Λ \Lambda Σ \Sigma  Ψ \Psi
                 Δ \Delta Ξ \Xi  Υ \Upsilon Ω \Omega
                 Θ \Theta Π \Pi  Φ \Phi

Table 4: Binary Relations.

You can produce corresponding negations by adding a \not command as prefix to the following symbols.

< <  > >  = =
≤ \leq or \le ≥ \geq or \ge ≡ \equiv
≪ \ll  ≫ \gg           ≐ \doteq
≺ \prec  ≻ \succ  ∼ \sim
⪯ \preceq  ⪰ \succeq  ≃ \simeq
⊂ \subset  ⊃ \supset  ≈ \approx
⊆ \subseteq  ⊇ \supseteq     ≅ \cong
⊏ \sqsubset ^{a} ⊐ \sqsupset ^{a} ⋈ \Join ^{a}
⊑ \sqsubseteq ⊒ \sqsupseteq ▷◁ \bowtie
∈ \in  ∋ \ni , \owns ∝ \propto
⊢ \vdash  ⊣ \dashv  ⊧ \models
| \mid  ∥ \parallel  ⊥ \perp
⌣ \smile  ⌢ \frown  ≍ \asymp
: :             ∉ \notin  ≠ \neq or \ne

aUse the latexsym package to access this symbol

Table 5: Binary Operators.

+ +  − -
± \pm  ∓ \mp  ◁ \triangleleft
· \cdot  ÷ \div  ▷ \triangleright
× \times  \ \setminus  ⋆ \star
∪ \cup  ∩ \cap  ∗ \ast
⊔ \sqcup  ⊓ \sqcap  ∘ \circ
∨ \vee , \lor  ∧ \wedge , \land  • \bullet
⊕ \oplus  ⊖ \ominus  ⋄ \diamond
⊙ \odot  ⊘ \oslash  ⊎ \uplus
⊗ \otimes  ◯ \bigcirc  ⨿ \amalg
△ \bigtriangleup ▽ \bigtriangledown † \dagger
⊲ \lhd ^{a}  ⊳ \rhd ^{a}  ‡ \ddagger
⊴ \unlhd ^{a}  ⊵ \unrhd ^{a}  ≀ \wr

Table 6: BIG Operators.

 ∑ \sum  ⋃ \bigcup  ⋁ \bigvee  ⨁ \bigoplus
 ∏ \prod  ⋂ \bigcap  ⋀ \bigwedge ⨂ \bigotimes
 ∐ \coprod ⨆ \bigsqcup  ⨀ \bigodot
 ∫ \int  ∮ \oint  ⨄ \biguplus

0.1 List of Mathematical Symbols  5

Table 7: Arrows.

← \leftarrow or \gets ⟵ \longleftarrow  ↑ \uparrow
→ \rightarrow or \to  ⟶ \longrightarrow  ↓ \downarrow
↔ \leftrightarrow  ⟷ \longleftrightarrow ↕ \updownarrow
⇐ \Leftarrow  ⟸ \Longleftarrow  ⇑ \Uparrow
⇒ \Rightarrow  ⟹ \Longrightarrow  ⇓ \Downarrow
⇔ \Leftrightarrow  ⟺ \Longleftrightarrow ⇕ \Updownarrow
↦ \mapsto  ⟼ \longmapsto  ↗ \nearrow
↩ \hookleftarrow  ↪ \hookrightarrow  ↘ \searrow
↼ \leftharpoonup  ⇀ \rightharpoonup  ↙ \swarrow
↽ \leftharpoondown  ⇁ \rightharpoondown  ↖ \nwarrow
⇌ \rightleftharpoons ⟺ \iff (bigger spaces)  ⇝ \leadsto ^{a}

aUse the latexsym package to access this symbol

Table 8: Delimiters.

 ( (  ) )  ↑ \uparrow  ⇑ \Uparrow
 [ [ or \lbrack ] ] or \rbrack ↓ \downarrow  ⇓ \Downarrow
 { \{ or \lbrace } \} or \rbrace ↕ \updownarrow ⇕ \Updownarrow
 ⟨ \langle  ⟩ \rangle  | | or \vert  ∥ \| or \Vert
 ⌊ \lfloor  ⌋ \rfloor  ⌈ \lceil  ⌉ \rceil

/ /  \ \backslash  . (dual. empty)

Table 9: Large Delimiters.

    { \lgroup     } \rgroup     { \lmoustache  } \rmoustache
    | \arrowvert  ‖ \Arrowvert  | \bracevert

\underline{6}

Table 10: Miscellaneous Symbols.

. . . \dots  · · · \cdots       ⋮ \vdots  ... \ddots
ℏ \hbar  ı \imath  ȷ \jmath  ℓ \ell
ℜ \Re  ℑ \Im  ℵ \aleph  ℘ \wp
∀ \forall  ∃ \exists  ℧ \mho ^{a}  ∂ \partial
′ '  ′ \prime  ∅ \emptyset ∞ \infty
∇ \nabla  △ \triangle □ \Box ^{a}  ◇ \Diamond ^{a}
⊥ \bot  ⊤ \top  ∠ \angle  √ \surd
♢ \diamondsuit ♡ \heartsuit ♣ \clubsuit ♠ \spadesuit
¬ \neg or \lnot ♭ \flat  ♮ \natural  ♯ \sharp

aUse the latexsym package to access this symbol

Table 11: Non-Mathematical Symbols.

These symbols can also be used in text mode.

                † \dag  § \S   \begin{matrix}c \\ ◯ \end{matrix} \copyright
                ‡ \ddag ¶ \P £ \pounds

Table 12: AMS Delimiters.

    ⌜ \ulcorner ⌝ \urcorner ⌞ \llcorner ⌟ \lrcorner

Table 13: AMS Greek and Hebrew.

 𝟋 \digamma κ \varkappa ℶ \beth ℸ \daleth ℷ \gimel

Table 14: AMS Binary Relations.

⋖ \lessdot  ⋗ \gtrdot  ≑ \doteqdot or \Doteq
⩽ \leqslant  ⩾ \geqslant  ≓ \risingdotseq
⪕ \eqslantless  ⪖ \eqslantgtr  ≒ \fallingdotseq
≦ \leqq  ≧ \geqq  ≖ \eqcirc
≪ \lll or \llless  ≫ \ggg or \gggtr  ≗ \circeq
≲ \lesssim  ≳ \gtrsim  ≜ \triangleq
⪅ \lessapprox  ⪆ \gtrapprox  ≏ \bumpeq
≶ \lessgtr  ≷ \gtrless  ≎ \Bumpeq
⋚ \lesseqgtr  ⋛ \gtreqless  ∼ \thicksim
⪋ \lesseqqgtr  ⪌ \gtreqqless  ≈ \thickapprox
≼ \preccurlyeq  ≽ \succcurlyeq  ≊ \approxeq
⋞ \curlyeqprec  ⋟ \curlyeqsucc  ∽ \backsim
≾ \precsim  ≿ \succsim  ⋍ \backsimeq
≾ \precapprox  ≿ \succapprox  ⊨ \vDash
⫅ \subseteqq  ⫆ \supseteqq  ⊩ \Vdash
⋐ \Subset  ⋑ \Supset  ⊪ \Vvdash
⊏ \sqsubset  ⊐ \sqsupset  ϶ \backepsilon
∴ \therefore  ∵ \because  ∝ \varpropto
 ∣ \shortmid  ∥ \shortparallel  ≬ \between
⌣ \smallsmile  ⌢ \smallfrown  ⋔ \pitchfork
◁ \vartriangleleft ▷ \vartriangleright ◀ \blacktriangleleft
⊴ \trianglelefteq  ⊵ \trianglerighteq  ▶ \blacktriangleright

Table 15: AMS Arrows.

←−− \dashleftarrow  −−→ \dashrightarrow  ⊸ \multimap
⇇ \leftleftarrows  ⇉ \rightrightarrows  ⇈ \upuparrows
⇆ \leftrightarrows  ⇄ \rightleftarrows  ⇊ \downdownarrows
⇚ \Lleftarrow  ⇛ \Rrightarrow  ↿ \upharpoonleft
↞ \twoheadleftarrow  ↠ \twoheadrightarrow ↾ \upharpoonright
↢ \leftarrowtail  ↣ \rightarrowtail  ⇃ \downharpoonleft
⇋ \leftrightharpoons ⇌ \rightleftharpoons ⇂ \downharpoonright
↰ \Lsh  ↱ \Rsh  ⇝ \rightsquigarrow
↫ \looparrowleft  ↬ \looparrowright  ↭ \leftrightsquigarrow
↶ \curvearrowleft  ↷ \curvearrowright
⟲ \circlearrowleft  ⟳ \circlearrowright

Table 16: AMS Negated Binary Relations and Arrows.

≮ \nless  ≯ \ngtr  ⫋ \varsubsetneqq
⪇ \lneq  ⪈ \gneq  ⫌ \varsupsetneqq
≰ \nleq  ≱ \ngeq  ⫅̸ \nsubseteqq
⩽̸ \nleqslant  ⩾̸ \ngeqslant  ⫆̸ \nsupseteqq
≨ \lneqq  ≩ \gneqq  ∤ \nmid
≨ \lvertneqq  ≩ \gvertneqq  ∦ \nparallel
≦̸ \nleqq  ≧̸ \ngeqq  ∤ \nshortmid
⋦ \lnsim  ⋧ \gnsim  ∦ \nshortparallel
⪉ \lnapprox  ⪊ \gnapprox  ≁ \nsim
⊀ \nprec  ⊁ \nsucc  ≇ \ncong
⪯̸ \npreceq  ⪰̸ \nsucceq  ⊬ \nvdash
⪵ \precneqq  ⪶ \succneqq  ⊭ \nvDash
⋨ \precnsim  ⋩ \succnsim  ⊮ \nVdash
⪹ \precnapprox  ⪺ \succnapprox  ⊯ \nVDash
⊊ \subsetneq  ⊋ \supsetneq  ⋪ \ntriangleleft
⊊ \varsubsetneq ⊋ \varsupsetneq ⋫ \ntriangleright
⊈ \nsubseteq  ⊉ \nsupseteq  ⋬ \ntrianglelefteq
⫋ \subsetneqq  ⫌ \supsetneqq  ⋭ \ntrianglerighteq
↚ \nleftarrow  ↛ \nrightarrow ↮ \nleftrightarrow
⇍ \nLeftarrow  ⇏ \nRightarrow ⇎ \nLeftrightarrow

Table 17: AMS Binary Operators.

∔ \dotplus  ▪ \centerdot  ⊺ \intercal
⋉ \ltimes  ⋊ \rtimes  ⋇ \divideontimes
⋓ \Cup or \doublecup ⋒ \Cap or \doublecap ∖ \smallsetminus
⊻ \veebar  ⊼ \barwedge  ⩞ \doublebarwedge
⊞ \boxplus  ⊟ \boxminus  ⊝ \circleddash
⊠ \boxtimes  ⊡ \boxdot  ⊚ \circledcirc
⋋ \leftthreetimes  ⋌ \rightthreetimes ⊛ \circledast
⋎ \curlyvee  ⋏ \curlywedge

Table 18: AMS Miscellaneous.

ℏ \hbar  ℏ \hslash  𝕜 \Bbbk
□ \square  ■ \blacksquare  Ⓢ \circledS
△ \vartriangle  ▲ \blacktriangle  ∁ \complement
▽ \triangledown ▼ \blacktriangledown ⅁ \Game
♢ \lozenge  ♦ \blacklozenge  ⋆ \bigstar
∠ \angle  ∡ \measuredangle  ∢ \sphericalangle
⧸ \diagup  ⧹ \diagdown  ‵ \backprime
∄ \nexists  Ⅎ \Finv  ∅ \varnothing
ð \eth  ℧ \mho

Table 19: Math Alphabets.

Example Command  Required package

ABCdef \mathrm{ABCdef}
ABCdef \mathit{ABCdef}
ABCdef \mathnormal{ABCdef}
ABC  \mathcal{ABC}
𝒜 ℬ𝒞  \mathcal{ABC}  mathrsfs

𝒜ℬ𝒞  \mathcal{ABC}  eucal with option: mathcal or
\mathscr{ABC}  eucal with option: mathscr

𝔄𝔅ℭ𝔡𝔢𝔣 \mathfrak{ABCdef}  eufrak

𝔸𝔹ℂ  \mathbb{ABC}  amsfonts or amssymb

0.1 List of Mathematical Symbols  11